# A Mathematical Framework for the Problem of Security for Cognition in Neurotechnology

### v1

@@ 5.2 Measurement Statistics
Because there already exists a robust body of literature on algorithms for classical HDC, including previously cited works such as,[15, 16, 18, 19, 37, 14] but also across the entire field of HDC, this represents a potentially very fruitful intersection and opportunity to adapt new algorithms to quantum computers – and perhaps to cognitive science as well. As an aside, while we anticipate that HDC’s usefulness in quantum computing will depend on context just as it has for classical computing, our hope is that it will not only bring new algorithms to quantum computing but also enable the development of novel quantum algorithms de novo. Reformulating information in terms of distances rather than specific values is a significant departure from the norm in quantum algorithms, but as we have shown, the robustness gains can be substantial. For Cognitive Security, however, the robustness of Projective Holographic Cognition is central.
@@ 5.3 Operational Dynamic Analysis
We now turn to discussing the practical implications of the statistical properties demonstrated above. Just as in Hyperdimensional Computing, favorable properties of high-dimensional geometry can be expected to make measurements of large numbers of cognitive states remarkably robust. Even with the significant degree of noise and myriad other challenges known to exist in many forms of psychological and cognitive assessment,[66, 67] HVs remain easily distinguishable on the scale of such high error rates as to at least mitigate most such effects. Further, correlational structure among cogits in a given PHV would enable error-correction. It is these correlations within individuals’ cogits, and those between the cogit HVs of different individuals, which are likely to enable greater robustness in machine learning models correlating neural activity with cognition as the scale of cogits considered increases. Additionally, the empirical fact that different people share beliefs in common to differing degrees shows that similarities in the correlation structures of different individuals’s cogit HVs are all but inevitable. This presents the opportunity for transfer learning from models derived from population-level experimentation over to models which target more specific groups – or perhaps even individuals.
@@
Similarly, models of the dynamics of cogit HVs over time will almost inevitably have correlations between individuals, because of the empirical fact that some people think more similarly than others. This extends the potential for transfer learning between different models linking neural activity with changes in cogits over time.
@@
Two main scenarios are of interest. The first is attempts by malicious actors to identify cognitive states against a subject’s will, and the second is attempts to modify cognitive states against the subject’s will. We term them Readout Attacks (RA) and Alteration Attacks (AA). These could range from blatantly dangerous situations like prisoners being interrogated by an authoritarian regime, to superficially benign yet still fundamentally malign cases like invasive methods of optimizing the manipulative efficacy of corporate advertising. Interrogation and focus groups, respectively, taken to new levels.
@@
In efforts to read out and/or modify cognitive activity, there are at least three broad levels at which machine learning models could be developed. Because we are not aware of any existing terminology for this topic, the three categories we anticipate are described below. The general goal of analyzing neural activity for cognitive readouts is to connect the dynamics at the relevant scale of monitoring with the cognitive operations they produce. We term this “Operational Dynamic Analysis”, or ODA. Of note, here “Dynamic” does not refer to the cogit dynamics, but instead to neural dynamics. Whether one is attempting to infer a cogit state HV at one approximate point in time or to infer the dynamics of cogit HVs over time, the dynamics of neural activity will be the data input into such models. The three subtypes we delineate are
@@
  • $\beta$-ODA : Methods of analysis which enable the learning and prediction of dynamical correlates of an individual’s cognition related to one or more topics, beliefs, cognitive processes, etc. via assessment across a group of similar individuals.
  • $\gamma$-ODA : Methods of analysis which enable the learning of precise dynamical correlates of one individual’s cognition related to specific topics, beliefs, cognitive processes, etc. by learning via assessment of a single individual.
There will presumably be different contexts where each proves most useful, but in general we imagine a progression of model generation and refinement which could move bidirectionally across the range from $\alpha$ to $\gamma$-ODA. Unless neural correlates of belief turn out to have such high variance between individuals that there is a complete lack of consistent patterns for transfer learning, or are so similar that refinement for individual subjects is unnecessary, one could benefit from constructing a general model from a population of subjects. Likewise, one could benefit in broadening or improving the efficacy of a population model via transfer learning from models developed from analyzing single individuals or groups of similar individuals. Given the inherent infeasibility of preventing malicious actors from accessing data from a population sample, [^27] State actors or other well-funded groups attempting such attacks would have access to the manpower and resources to collect their own data and build their own models. and all the more so data on small groups or individuals, security considerations will need to center on hampering attacks at the level of individual brains. Acknowledging the impossibility of preventing malicious actors from developing tools, we instead focus on the need for strategic defense methods.
@@ 6 Readout Attacks
A general means of interfering with RA would be to inject noise, much as with the reconstruction attacks of Differential Privacy.[12, 13] The goal would be to inject noise such that a metric of model dissimilarity $S$ is maximized over minimal noise. For the sake of ease of analysis, we will assume it is technological in origin to improve flexibility in optimizing noise $\aleph$.
@@ 6.1 Model-Aware Defense
We will begin by discussing an approach which assumes some level of knowledge of the possible models attackers may leverage. In a later section we will discuss defense which does not assume any knowledge of the potential models or model types which attackers may employ. However, any bounds on such a Model-Agnostic Defense would almost certainly be sub-optimal for many possible models. To assess Model-Aware Defense one would ideally like to be able to define some analogue of a cryptographic universe. However, the complexity of the Cognitive Security problem and breadth of potential elements of $H_{A}$ and $H_{D}$ makes constructing such universally strong theorems challenging – if not impossible. Instead, we acknowledge the limitations of Model-Aware Defense and present relevant ideas before moving to their generalizations in Model-Agnostic Defense.
@@
The defenders infer $\eta(\aleph)$ from assessment of how different $\aleph$ impact cognition. The quantum JS divergence (QJSD) is chosen because of its normalization, given the Kullback-Liebler divergence (KLD) is unbounded. We multiply the cognitive noise $\eta$ by $-1$ so that the $D_{QJS}$ quantifies the information needed to identify $A$, as the JSD and its quantum equivalent measure the information needed to identify a mixture of distributions given one of the distributions. In this case, given post-noise observations $B$ we want to measure the information needed to identify $A$, which is $B-\eta$. With the $-\lambda$ weighting, we seek to minimize this additional information. The goal of an attacker will be to identify a true distribution given observations from some altered distribution, the inefficiency of which is measured by the JSD (and in this case, its quantum analog). In this manner, one aims to leverage the intrinsic noisiness and robustness of biological neural computing in order to confound readouts by producing properties which maximally exploit out-of-distribution weaknesses of readout models, while preserving the underlying cognition itself.
@@ 6.2 Model-Agnostic Defense
Any system for an attacker to read out cognitive states will need to use information-theoretic properties of brain activity in order to correlate them with cognition. This is the space in which $\epsilon$ operates. Meanwhile, we would like to perturb cognitive state $\psi$ as little as possible. This allows us to define optimal noise in a way which is more fundamental, though we do not make conjectures as to whether it or State Model-Optimal Noise (SMON)/Dynamics Model-Optimal Noise (DMON) will be more useful in any given context.
@@
Let $D_{JS}$ be the classical Jensen-Shannon divergence and $D_{QJS}$ the quantum Jensen-Shannon divergence. Assuming a projective cognition model and a classical description of neural activity, for baseline neural state distribution $Q$, neural noise distribution $\aleph$, empiric cognitive noise distribution $\eta(\aleph)$, baseline cognitive distribution $A$, and post-noise cognitive distribution $B=A+\eta$, distribution $\aleph$ is state information-optimal if for a constraining parameter $\lambda$ and a required minimum dissimilarity $\mu\in[0,1]$,
$\hat{\aleph}=$ $\arg\max_{\aleph}\left(D_{JS}(Q,-\aleph)-\lambda D_{QJS}(B,-\eta(\aleph))\right)$
@@
$D_{JS}(Q,Q+\aleph)\geq\mu\hskip 28.45274pt\forall F\in H$
@@
Let $D_{JS}$ be the classical Jensen-Shannon divergence and $D_{QJS}$ the quantum Jensen-Shannon divergence. Assuming a projective cognition model and a classical description of neural activity, for baseline neural state distribution $Q$, neural noise distribution $\aleph$, function $\kappa(\aleph)$ mapping neural activity noise to cognitive dynamics noise, baseline cognitive dynamics operator distribution $\Lambda$, and post-noise cognitive distribution $\Delta=\Lambda+\kappa(\aleph)$, distribution $\aleph$ is state information-optimal if for a constraining parameter $\lambda$ and a required minimum dissimilarity $\mu\in[0,1]$,
$\hat{\aleph}=$ $\arg\max_{\aleph}\left(D_{JS}(Q,-\aleph)-\lambda D_{QJS}(\Delta,-\kappa(\aleph))\right)$
@@ 7 Alteration Attacks
With $D_{JS}$ the Jensen-Shannon divergence, $D_{QJS}$ the quantum Jensen-Shannon divergence, let $F$ a model mapping from distributions of neural activity to distributions of cognitive states, $Y(\cdot,B)$ a normalized objective function of how close some distribution is to a target cognitive state distribution $B$, and $\aleph$ the distribution of an input signal altering neural activity. For baseline neural state distribution $Q$, baseline cognitive distribution $A\approx F(Q)$, and desired post-alteration cognitive distribution $C\approx F(Q+\aleph)$, alteration signal $\aleph$ is model-alteration optimal if for a constraining parameter $\lambda$
$\hat{\kappa}=$ $\arg\max_{\aleph}Y(F(Q+\aleph),B)-\lambda D_{JS}(Q,Q+\aleph)$
This definition is in some sense an inversion of that used by defenders. An attacker seeks to minimize the difference between the distributions of neural states, described by the JS-divergence, to avoid detection. Simultaneously, they aim to minimize the difference between the target cognitive state distribution $B$ and the cognitive state distribution inferred via the attacker’s model, $F(Q+\aleph)$. In analogy with the defenders’ own noise signal, here $\aleph$ is the targeted signal applied by attackers.
@@
With $D_{JS}$ the Jensen-Shannon divergence, $D_{QJS}$ the quantum Jensen-Shannon divergence, let $G$ a model mapping from distributions of neural activity to distributions of cognitive states, $\Phi$ a target distribution of cognitive dynamics operators, and $\aleph$ the distribution of an input signal altering neural activity. For baseline neural state distribution $Q$, distribution $\aleph$ is dynamics model-optimal if for a constraining parameter $\lambda$
$\hat{\aleph}=$ $\arg\max_{\aleph}\left(\min_{G\in H}\hskip 2.84526pt\frac{1}{2}\mathbf{E}(||G(Q+\aleph)-\Phi||)\right)-\lambda D_{JS}(Q,Q+\aleph)$
These two definitions describe the attackers’ efforts to optimize the tradeoff between 1) how closely they approach the desired effect and 2) how readily a defender could detect this alteration attack. The former is described by the first term in the objective functions, and the latter described by $D_{JS}(Q,Q+\aleph)$. Minimizing this divergence term reduces the possible information available for defenders to detect.
@@ 8 Conclusions
In this paper we aimed to initiate an entirely new direction of research. By linking Projective (“Quantum”) Cognition – the unorthodox yet highly effective modeling approach which resolves the major pardoxes of set-theoretic models of cognition – with ideas from quantum computing, Hyperdimensional Computing, and Differential Privacy, we have provided a framework for precisely analyzing issues related to the fast-emerging problem of Cognitive Security. This represents a novel intersection between research on cognitive science, machine learning, privacy and security, artificial intelligence, and neurotechnology, and one which we expect will have significant practical importance in the coming years. The exponential progress of neurotechnology is enabling ever greater medical advancements, but as with all tools it has the potential for both benevolent and malign uses. To the best of our knowledge all existing technical literature has focused on the former, but as has been made clear by the history of computer and data security, if a tool can be used to violate security and privacy for the gain of malicious actors, it will be used for precisely that purpose. There is no reason to assume neurotechnology will be any exception, and such a gamble would pose great risk.
@@
Additionally, there are a few interesting notions which arise from the links between fields which have been forged in this paper. First, that quantum computers could be used for simulating cognitive dynamics. Second, that it may be feasible to develop quantum algorithms or AI systems which interface with human intuition in new, more natural ways than the available classical systems do. Separate from these questions, it is interesting to note that there is nothing requiring the Cognitive Security framework be restricted to the individual level. This is not in the same sense as the $\alpha$/$\beta$/$\gamma$-ODA transfer learning described in Section 5.3, but in a more general way. Cognitive Security is deliberately a general framework to enable greater flexibility, and we suspect it may also prove effective in a somewhat scale-agnostic manner to describe the cognitive dynamics of groups of people – given the right modifications to the above objective functions. The potential implications and applications of all three possibilities are, however, much beyond the scope of this paper.
@@
The question has justifiably been raised of whether it is ethical and safe to put forward the preceding descriptions of the structures of different potential attack types. [^31] Personal communication, Professor Milana Boukhman Trounce, Director of Biosecurity at Stanford University and founding Chair of Biosecurity at the American College of Emergency Physicians. Indeed, it is likely that some malicious actors will work to leverage these ideas for unethical purposes. But supposing they were not made public, practically speaking it would be impossible to selectively communicate them solely to parties who would use them only for the common good. And in such a case, any attempts at secrecy would be more likely to mainly impact some portion of those who would use it for benevolent ends than those who would use it for malicious ones. As such, we consider the most ethical option to be the release of this paper to the public. While Cognitive Security – being a new area – is not a traditional domain of Biosecurity, some experts have argued that it will not only very soon become one, but will be a radically new world of opportunities and challenges. Such opportunities and challenges are likely to include some which are analogous to biosecurity problems of other modalities, and many which are wholly unlike anything seen before.[68] This is a view we vehemently share.
@@
Finally, there is an interesting notion which emerges from the discussion of attack and defense methods. Cognitive Security against readout requires alteration, albeit strategically minimal alteration. And as seen with the similarities between SMON/SION and MOA/IOA, there is a meaningful analogy between AA and defensive measures against RA. If no readout is made of cognition and neural activity, then alteration may not in general be reliably detected. The roles of the attacker and the defender exist as odd sorts of mirror images in both cases. While we do not currently see any path towards proving this seeming principle one way or another, it leads us to make the following conjecture. There may be no means to achieve Cognitive Security measures against readout without enacting some degree of alteration to cognition itself, nor any efficient measures against alteration without performing some readout. As with much of the myriad implications of this paper and the problems it seeks to begin addressing, there is a great deal of future work to be done on the topic.

### v4

@@ 3.2 Measurement Statistics
Because there already exists a robust body of literature on algorithms for classical HDC, including previously cited works such as,[klyeko_survey_hyperdimensional_computing_part_I, klyeko_survey_hyperdimensional_computing_part_II, Neubert_2019_hyperdimensional_computing_for_robotics, Karunaratne_2020_electronics_with_hyperdimensional_computing, kim_hyper_dimensional_complex_vector_low-power_communication, Kanerva_2009_introduction_to_distributed_representations_hyperdimensional_computing] but also across the entire field of HDC, this represents a potentially very fruitful intersection and opportunity to adapt new algorithms to quantum computers – and perhaps to cognitive science as well. As an aside, while we anticipate that HDC’s usefulness in quantum computing will depend on context just as it has for classical computing, our hope is that it will not only bring new algorithms to quantum computing but also enable the development of novel quantum algorithms de novo. Reformulating information in terms of distances rather than specific values is a significant departure from the norm in quantum algorithms, but as we have shown, the robustness gains can be substantial. For Cognitive Neurosecurity, however, the robustness of Projective Holographic Cognition is central.
@@ 3.3 Operational Dynamic Analysis
We now turn to discussing the practical implications of the statistical properties demonstrated above. Just as in Hyperdimensional Computing, favorable properties of high-dimensional geometry can be expected to make measurements of large numbers of cognitive states remarkably robust. Even with the significant degree of noise and myriad other challenges known to exist in many forms of psychological and cognitive assessment,[Green_2011_psych_testing_assessment_reliability, Meyer_2001_psych_testing_and_assessment_review_of_issues] hypervectors remain easily distinguishable on the scale of such high error rates as to at least mitigate most such effects. Further, correlational structure among cogits in a given projective hypervector would enable error-correction. It is these correlations within individuals’ cogits, and those between the cogit hypervectors of different individuals, which are likely to enable greater robustness in machine learning models correlating neural activity with cognition as the scale of cogits considered increases. Additionally, the empirical fact that different people share beliefs in common to differing degrees shows that similarities in the correlation structures of different individuals’s cogit hypervectors are all but inevitable. This presents the opportunity for transfer learning from models derived from population-level experimentation over to models which target more specific groups – or perhaps even individuals.
@@
Similarly, models of the dynamics of cogit hypervectors over time will almost inevitably have correlations between individuals, because of the empirical fact that some people think more similarly than others. This extends the potential for transfer learning between different models linking neural activity with changes in cogits over time.
@@
Two general brainjacking scenarios are of interest. The first is attempts by malicious actors to identify cognitive states against a subject’s will, and the second is attempts to modify cognitive states against the subject’s will. We term them Readout Attacks (RA) and Alteration Attacks (AA). These could range from blatantly dangerous situations like prisoners being interrogated by an authoritarian regime, to superficially benign yet still fundamentally malign cases like invasive methods of optimizing the manipulative efficacy of advertising or propaganda. Interrogation and focus groups, respectively, taken to new levels. A fine-grained survey of attack types is given by Bernal et al,[bernal_2022_bci_security_review] but for the purposes of this paper we will focus on these two broad categories rather than expanding the discuss to the fine-grained level of differing attack types provided in their excellent overview. [^18] Essentially all items on the Bernal et al list can be collected under these two broad categories, so follow-up work to this paper could explore with greater depth the application of this paper’s framework to the subtypes they present.
@@
In efforts to read out and/or modify cognitive activity, there are at least three broad levels at which machine learning models could be developed. Because we are not aware of any existing terminology for this topic, the three categories we anticipate are described below. The general goal of analyzing neural activity for cognitive readouts is to connect the dynamics at the relevant scale of monitoring with the cognitive operations they produce. We term this “Operational Dynamic Analysis”, or ODA. The three subtypes we delineate are
@@
  • $\beta$-ODA : Methods of analysis which enable the learning and prediction of dynamical correlates of an individual’s cognition related to one or more topics, beliefs, cognitive processes, etc. via assessment across a group of individuals who are similar according to chosen metric(s) of interest.
  • $\gamma$-ODA : Methods of analysis which enable the prediction of dynamical correlates of one individual’s cognition related to specific topics, beliefs, cognitive processes, etc. by learning via assessment of a single individual.
There will presumably be different contexts where each proves most useful, but in general we imagine a progression of model generation and refinement which could move bidirectionally across the range from $\alpha$ to $\gamma$-ODA. Unless neural correlates of belief turn out to have such high variance between individuals that there is a complete lack of consistent patterns for transfer learning, or are so similar that refinement for individual subjects is unnecessary, one could benefit from constructing a general model from a population of subjects. In fact, recent work has shown this exact approach offers significant improvements for the well-studied task of reading out mental states from fMRI scans.[scotti_2024_transfer_learning_MRI_brain_readout] Likewise, one could benefit in broadening or improving the efficacy of a population model via transfer learning from models developed from analyzing single individuals or groups of similar individuals. Given the inherent infeasibility of preventing malicious actors from accessing data from a population sample, [^19] State actors or other well-funded groups attempting such attacks would have access to the manpower and resources to collect their own data and build their own models. and all the more so data on small groups or individuals, security considerations will need to center on hampering attacks at the level of individual brains. Acknowledging the impossibility of preventing malicious actors from developing tools, we instead focus on the need for strategic defense methods.
@@ 4 Readout Attacks
A general means of interfering with RA would be to inject noise, much as with the reconstruction attacks of Differential Privacy.[Dwork_roth_2014_differential_privacy_textbook, wood_2020_differential_privacy_for_non_technical_audience] This can be done via the Bonaci et al BCI Anonymizer,[bonaci_2014_brain_security_exocortex] assuming no air-gap attacks are developed for current and future BCI technologies. However, one could also use a secondary implanted system which introduces noise into the neuronal function itself. [^20] Such a secondary device would need to be physically separate from the main device to at least some extent in order to mitigate air-gap attacks, as magnetic and acoustic air-gap attacks would likely be difficult to defend against without such separation. In this case, the brain’s own magnetic activity would serve to shield a physically distanced noise generator from potential air-gapped magnetic attack via the main interface. Given a sufficiently complex time-varying function for generating a noise kernel and experimentation to optimize such hardware and the noise signals they produce, a physically hardened device would be resistant to both traditional and air-gap attacks while providing a form of privacy much more difficult to circumvent. The goal would be to inject noise such that a metric of model dissimilarity $S$ is maximized over noise which minimizes impact on cognition. For the sake of ease of analysis, we will assume it is technological in origin to improve flexibility in optimizing noise $\aleph$. While we envisage this as described above, the mathematical statements apply equally to something akin to the BCI Anonymizer,[bonaci_2014_brain_security_exocortex] replacing induced physiologic noise with noise injected between the recording portion of the hardware and the portion performing computations. For AA there would in fact be no difference if it were physiologic noise vs. something like a BCI Anonymizer, but for RA the below would become more similar to the Bonaci et al approach. Our results thus remain applicable even for more indirect approaches to security, in the vein of BCI hardware modifications. However, such approaches retain the greater vulnerability to potential air-gap attacks, and given the stakes and potential gain for malicious actors we expect a great deal of thought and malign innovation will be directed towards developing such attacks as BCI become more common.
@@ 4.1 Model-Aware Defense
We will begin by discussing an approach which assumes some level of knowledge of the possible models attackers may leverage. In a later section we will discuss defense which does not assume any knowledge of the potential models or model types which attackers may employ. However, any bounds on such a Model-Agnostic Defense would almost certainly be sub-optimal for many possible models. To assess Model-Aware Defense one would ideally like to be able to define some analogue of a cryptographic universe. However, the complexity of the Cognitive Neurosecurity problem and breadth of potential elements of $H_{A}$ and $H_{D}$ makes constructing such universally strong theorems challenging – if not impossible. Instead, we acknowledge the limitations of Model-Aware Defense and present relevant ideas before moving to their generalizations in Model-Agnostic Defense.
@@
The defenders infer $\eta(\aleph)$ from assessment of how different $\aleph$ impact cognition. The quantum Jensen-Shannon divergence (QJSD) is chosen because of its normalization, given the Kullback-Liebler divergence (KLD) is unbounded. We multiply the cognitive noise $\eta$ by $-1$ so that the $D_{QJS}$ quantifies the information needed to identify $A$, as the JSD and its quantum equivalent measure the information needed to identify a mixture of distributions given one of the distributions. In this case, given post-noise observations are drawn from the mixture distribution $B=\frac{1}{2}(A+\eta)$ we want to measure the information needed to identify $A$. This is precisely what the JSD and QJSD measure, as it describes the mutual information between a random variable sampled from a mixture distribution – true cognitive function plus noise – and an indicator variable denoting which of the two component distributions produced a given value – whether it is protective noise or underlying ground truth. With the $-\lambda$ weighting, we seek to minimize this additional information. The goal of an attacker will be to identify a true distribution given observations from some altered distribution, the inefficiency of which is measured by the JSD (and in this case, its quantum analog). In this manner, one aims to leverage the intrinsic noisiness and robustness of biological neural computing in order to confound readouts by producing properties which maximally exploit out-of-distribution weaknesses of readout models, while preserving the underlying cognition itself.
@@ 4.2 Model-Agnostic Defense
Any system for an attacker to read out cognitive states will need to use information-theoretic properties of brain activity in order to correlate them with cognition. As laid out very neatly by Bonaci et al,[bonaci_2014_brain_security_exocortex] such information-theoretic descriptions provide a strong foundation for problems of this sort. Here we adapts and extend such ideas so as to interface with the precise mathematical description of cognitive states. This is the space in which $\epsilon$ operates. Meanwhile, we would like to perturb cognitive state $\psi$ as little as possible. This allows us to define optimal noise in a way which is more fundamental, though we do not make conjectures as to whether it or State Model-Optimal Noise (SMON)/Dynamics Model-Optimal Noise (DMON) will be more useful in any given context as there are tradeoffs for each.
@@
Let $D_{J}$ be the classical Jensen-Shannon divergence and $D_{QJS}$ the quantum Jensen-Shannon divergence. Assuming a projective cognition model and a classical description of neural activity, for baseline neural state distribution $Q$, neural noise distribution $\aleph$, empiric cognitive noise distribution $\eta(\aleph)$, baseline cognitive distribution $A$, and post-noise cognitive distribution $B=A+\eta$, distribution $\aleph$ is state information-optimal if for a constraining parameter $\lambda$ and a required minimum dissimilarity $\mu\in[0,1]$,
$\hat{\aleph}=$ $\arg\max_{\aleph}\left(D_{J}(Q,-\aleph)-\lambda D_{QJS}(B,-\eta(\aleph))\right)$
@@
$D_{J}(Q,Q+\aleph)\geq\mu\hskip 28.45274pt\forall F\in H$
@@
Let $D_{J}$ be the classical Jensen-Shannon divergence and $D_{QJS}$ the quantum Jensen-Shannon divergence. Assuming a projective cognition model and a classical description of neural activity, for baseline neural state distribution $Q$, neural noise distribution $\aleph$, function $\kappa(\aleph)$ mapping neural activity noise to cognitive dynamics noise, baseline cognitive dynamics operator distribution $\Lambda$, and post-noise cognitive distribution $\Delta=\Lambda+\kappa(\aleph)$, distribution $\aleph$ is state information-optimal if for a constraining parameter $\lambda$ and a required minimum dissimilarity $\mu\in[0,1]$,
$\hat{\aleph}=$ $\arg\max_{\aleph}\left(D_{J}(Q,-\aleph)-\lambda D_{QJS}(\Delta,-\kappa(\aleph))\right)$
@@ 5 Alteration Attacks
With $D_{J}$ the Jensen-Shannon divergence, $D_{QJS}$ the quantum Jensen-Shannon divergence, let $F$ a model mapping from distributions of neural activity to distributions of cognitive states, $Y(\cdot,B)$ a normalized objective function of how close some distribution is to a target cognitive state distribution $B$, and $\aleph$ the distribution of an input signal altering neural activity. For baseline neural state distribution $Q$, baseline cognitive distribution $A\approx F(Q)$, and desired post-alteration cognitive distribution $C\approx F(Q+\aleph)$, alteration signal $\aleph$ is model-alteration optimal if for a constraining parameter $\lambda$
$\hat{\kappa}=$ $\arg\max_{\aleph}Y(F(Q+\aleph),B)-\lambda D_{J}(Q,Q+\aleph)$
This definition is in some sense an inversion of that used by defenders. An attacker seeks to minimize the difference between the distributions of neural states, described by the J-divergence, to avoid detection. Simultaneously, they aim to minimize the difference between the target cognitive state distribution $B$ and the cognitive state distribution inferred via the attacker’s model, $F(Q+\aleph)$. In analogy with the defenders’ own noise signal, here $\aleph$ is the targeted signal applied by attackers.
@@
With $D_{J}$ the Jensen-Shannon divergence, $D_{QJS}$ the quantum Jensen-Shannon divergence, let $G$ a model mapping from distributions of neural activity to distributions of cognitive states, $\Phi$ a target distribution of cognitive dynamics operators, and $\aleph$ the distribution of an input signal altering neural activity. For baseline neural state distribution $Q$, distribution $\aleph$ is dynamics model-optimal if for a constraining parameter $\lambda$
$\hat{\aleph}=$ $\arg\max_{\aleph}\left(\min_{G\in H}\hskip 2.84526pt\frac{1}{2}\mathbf{E}(||G(Q+\aleph)-\Phi||)\right)-\lambda D_{J}(Q,Q+\aleph)$
These two definitions describe the attackers’ efforts to optimize the tradeoff between 1) how closely they approach the desired effect and 2) how readily a defender could detect this alteration attack. The former is described by the first term in the objective functions, and the latter described by $D_{J}(Q,Q+\aleph)$. Minimizing this divergence term reduces the possible information available for defenders to detect.
@@ 6 Conclusions
In this paper we aimed to initiate an entirely new direction of research. By linking Projective (“Quantum”) Cognition – the unorthodox yet highly effective modeling approach which resolves the major pardoxes of set-theoretic models of cognition – with ideas from quantum computing, Hyperdimensional Computing, and Differential Privacy, we have provided a framework for precisely analyzing issues related to the fast-emerging problem of Cognitive Neurosecurity. This represents a novel intersection between research on cognitive science, machine learning, privacy and security, artificial intelligence, and neurotechnology, and one which we expect will have significant practical importance in the coming years. The exponential progress of neurotechnology is enabling ever greater medical advancements, but as with all tools it has the potential for both benevolent and malign uses. To the best of our knowledge all existing technical literature has focused on the former, but as has been made clear by the history of computer and data security, if a tool can be used to violate security and privacy for the gain of malicious actors, it will be used for precisely that purpose. There is no reason to assume neurotechnology will be any exception, and such a gamble would pose great risk.
@@
Additionally, there are a few interesting notions which arise from the links between fields which have been forged in this paper. First, that quantum computers could be used for simulating cognitive dynamics at scale. This is of particular interest as some devices have reached scales relevant to holographic cognition.[Atom_Computing_2023, Ma_2023_atom_computing_1, Evered_2023_atom_computing_2, Scholl_2023_atom_computing_3] Second, that it may be feasible to develop quantum algorithms or AI systems which interface with human intuition in new, more natural ways than the available classical systems do. Separate from these questions, it is interesting to note that there is nothing requiring the Cognitive Neurosecurity framework be restricted to the individual level. This is not in the same sense as the $\alpha$/$\beta$/$\gamma$-ODA transfer learning described in Section 3.3, but in a more general way. Cognitive Neurosecurity is deliberately a general framework to enable greater flexibility, and we suspect it may also prove effective in a somewhat scale-agnostic manner to describe the cognitive dynamics of groups of people – given the right modifications to the above objective functions. Should BCIs become prevalent beyond the cases of medical necessity for which they are currently used, such coordinated attacks on groups will become increasingly viable and dangerous. Even in the absence of widespread use of neurotechnology, correlates besides neural activity could readily be used for the algorithmic problems described in this paper. In this regard, one could apply the CNS framework to existing forms of cognitive readout such as social media, with the potential advantage over existing methods being that holographic cognition is built on a first-principles empirical description of human cognition. The potential implications and applications of all these possibilities are, however, much beyond the scope of this paper. Additional follow-up work could apply this framework in more specific ways to each of the relevant attack types described in the Bernal et al review.[bernal_2022_bci_security_review]
@@
The question has justifiably been raised of whether it is ethical and safe to put forward the preceding descriptions of the structures of different potential attack types given they also give a mathematical description of a means of accomplishing a broader range of attacks than previously described in the literature. [^24] Personal communication, Professor Milana Boukhman Trounce, Director of Biosecurity at Stanford University and founding Chair of Biosecurity at the American College of Emergency Physicians.[bernal_2022_bci_security_review] This is particularly true due to the scale-agnostic nature of our approach. Indeed, it is likely that some malicious actors will work to leverage these ideas for unethical purposes. But supposing they were not made public, practically speaking it would be impossible to selectively communicate them solely to parties who would use them only for the common good. And in such a case, any attempts at secrecy would be more likely to mainly impact some portion of those who would use it for benevolent ends than those who would use it for malicious ones. As such, we consider the most ethical option to be the release of this paper to the public. While Cognitive Neurosecurity – being a new area – is not a traditional domain of Biosecurity, some experts have argued that it will not only very soon become one, but will be a radically new world of opportunities and challenges. Such opportunities and challenges are likely to include some which are analogous to biosecurity problems of other modalities, and many which are wholly unlike anything seen before.[poste_biosecurity_talk_stanford] This is a view we vehemently share.
@@
Finally, there is an interesting notion which emerges from the discussion of attack and defense methods. The Bonaci et al BCI Anonymizer is an innovative approach to the proper design of BCI hardware, but remains vulnerable to air-gap attacks, malicious implants, or BCI hardware vulnerabilities. In order to maximally mitigate such vulnerabilities, one would likely need to induce physiologic noise. This tactic for Cognitive Neurosecurity against readout requires alteration, albeit strategically minimal alteration. And as seen with the similarities between SMON/SION and MOA/IOA, there is a meaningful analogy between AA and defensive measures against RA, and between RA and defensive measures against AA. If no readout is made of cognition and neural activity, then alteration may not in general be reliably detected. The roles of the attacker and the defender exist as odd sorts of mirror images in both cases. While we do not currently see any path towards verifying this one way or another, it leads us to make the following conjecture. There may be no means to achieve physiology-level security guarantees against readout without enacting some degree of alteration to cognition itself, nor such guarantees against alteration without performing some readout. As with much of the myriad implications of this paper and the problems it seeks to begin addressing, there is a great deal of future work to be done on the topic.